# Beyond Walking: A Large-Scale Image-Text Benchmark for Text-based Person Anomaly Search

### v1

@@ 3.3 Dataset Analysis
Following the steps above, we have successfully created the Pedestrian Anomaly Behavior (PAB) dataset, a large-scale, richly annotated collection. Its statistics are detailed in Table 2, with properties and examples depicted in Figure 2. Table 1 compares PAB with other prominent Text-based Person Re-Identification datasets, such as CUHK-PEDES [24], ICFG-PEDES [11], and RSTPReid [55], focusing on image count and size, text annotation quantity and length, data sources, and annotation types. PAB features the following characteristics: (1) A Large Number of Anomalous Behaviors: Unlike general pedestrian datasets that focus on the viewpoint and occlusion of pedestrian images, PAB emphasizes providing a large number of images and textual descriptions of pedestrian anomaly. This complements ordinary Person Re-Identification tasks and presents new challenges. (2) High-Fidelity Images: Compared to pedestrian images from surveillance cameras, which often suffer from poor lighting, blurry textures, the images in PAB are of higher quality due to the generation method we adopted. The synthesized images are reasonable, realistic, and aesthetically pleasing. (3) Specific Textual Descriptions: Relative to existing cross-modal pedestrian datasets, the texts in PAB are longer and provide more detailed information, including the appearance, clothing, actions, and background context of individuals. For pedestrian anomaly retrieval, information about the appearance and actions of individuals is crucial. (4) Diversity: PAB encompasses a wide range of images that vary in terms of appearance, posture, viewpoint, background, and occlusion. Additionally, the text generation approach ensures that the image captions in PAB are sufficiently diverse, too. (5) Large-Scale Image-Text Pairs: PAB comprises over one million image-text pairs, supporting deep cross-modal models in learning better uni-modal features and more inter-modal associations from the data. (6) Rich Annotations: Each image-text pair in PAB is annotated with corresponding action, anomaly, and scene categories, providing additional data information besides person appearance. (7) Less Privacy Concerns: Apart from a small amount of data in the test set sourced from existing public datasets, the vast majority of the data in PAB is synthesized, reducing ethical and legal issues.
@@ 4.1 Pose-aware Image Encoder (PE)
Recognizing that normal and anomaly actions notably differ in human pose, we devise a pose-aware image encoder that explicitly incorporates human pose to enhance behavior comprehension. For a given image $I$, a human key point detector [7] is utilized to obtain the related pose map $P$. $I$ and $P$ are separately fed into the Image Encoder to efficiently generate image embedding $f_{I}$ and pose embedding $f_{P}$.
Image Encoder. Without loss of generality, we deploy Swin Transformer (Swin-B) [29] as the Image Encoder. The input image ($I$ or $P$) is initially divided into non-overlapping patches, which are then linearly embedded. These embedded patches are subsequently processed by transformer blocks, comprising Multi-Mead Self-Attention and Feed Forward modules, to generate patch embeddings. Each patch embedding encapsulates the information of its corresponding patch. To aggregate the information from all patches, we compute the average of their features, referred to as the [CLS] embedding, and prepend it to the sequence.
Pose-aware Cross-Attention. Pose representation $f_{P}$, after being regularized by Layer Norm (LN), is integrated into the image representation through a multi-head Cross-Attention module (CA). The output ($f_{CA}$) of the Cross-Attention can be defined by the following equation:
$f_{CA}$ $=\text{CA}({\bf q,k,v})=Softmax(\frac{{\bf qk}^{T}}{\sqrt{d}}){\bf v},$ (1)
${\bf q}$ $=W_{q}f_{P},\quad{\bf k}=W_{k}f_{I},\quad{\bf v}=W_{v}f_{I},$ (2)
where $q,k,v$ are the query, key, and values matrices of the attention operation respectively, and $W_{q},W_{k},W_{v}$ are the weight matrices of the trainable linear projection layers. Then, we simply add the output of CA to the image embedding $f_{I}$. The new image representation of $I$, terms as $f_{V}=f_{I}+f_{CA}$.
@@ 4.2 Identity-based Hard Negative Mining (IHNM)
The key to solving pedestrian anomaly research is effectively establishing the relationship between pedestrian descriptions (with normal or anomaly action) and images. Ideally, for each text, the model should be able to search for an image that matches the appearance, actions, and background described, particularly the action description. Training such a cross-modal search model requires large-scale matching image-text pairs, negative pairs, and hard negative pairs, which are crucial for enabling the model to learn more discriminative features. Based on the construction process of PAB, we propose an Identity-based hard negative pair mining (IHNM) method that provides corresponding hard negative samples for each image and text in the training set. The sampling strategy is illustrated in Figure 3 (b).
Consider a training pair $(I,T)$, where $I$ is generated from anomaly/normal caption $C\in\{C_{a},C_{a+},C_{n}\}$, and $T$ is its re-captioned text. As the normal/anomaly counterpart $\tilde{C}$ of $C$ (both describing the same person) is known according to the OOPS! dataset, we can pick $\tilde{I}$ generated from $\tilde{C}$, and its re-captioned text $\tilde{T}$ can also be chosen accordingly. As shown in the Figure 3 (b), $C$ and $\tilde{C}$ describe the anomalous and normal behaviors of the same pedestrian, respectively. Consequently, the generated images have similar pedestrian appearances and backgrounds but different actions. Intuitively, the image $I$ and the text $\tilde{T}$ form an ID-based hard negative pair. The establishment of hard negative pairs with similar appearances and backgrounds but different actions is critical for the model to learn features that are more discriminative of actions.
Method #Training IHNM PE R@1 R@5 R@10 mAP
Images
Zero-Shot 9.40 22.14 30.18 16.37
Baseline 0.1M 69.92 95.60 97.32 81.53
M1 0.1M $\checkmark$ 72.25 95.91 98.03 83.07
M2 0.1M $\checkmark$ 71.79 95.40 97.83 82.79
CMP 0.1M $\checkmark$ $\checkmark$ 72.80 96.01 97.47 83.47
(Ours) 1M $\checkmark$ $\checkmark$ 79.53 97.93 98.84 88.17
(a) Quantitative results and ablations on the key components of our method.
(b) Ablation studies on synthetic training data scale.
Figure 4: Quantitative results and Ablation studies of Text-based Person Anomaly Search on PAB. (a) Quantitative results of Text-based Person Anomaly Search on PAB and Ablation studies on the key component of our proposed method, i.e., Identity-based Hard Negative Mining (IHNM) and Pose-aware Image Encoder (PE). (b) We apply $0,10\%,25\%,50\%,100\%,$ data pairs of the PAB training set to train and then report the recall rate on the real-world test set.
@@ 4.3 Cross-modal Modeling
IHNM-Enhanced Image-Text Matching. Image-text matching determines whether an (image, text) pair is matched or not. Typically, well-aligned pairs are considered positive samples. Rather than randomly selecting trivial negatives, we employ our IHNM strategy to identify hard negative pairs. These pairs differ only in the action with its positive counterpart, requiring the model to discern subtle action differences, a capability crucial for the task of person anomaly search. Given an image-text pair, their multi-modal feature is encoded by feeding their representation into a cross encoder. Specifically, we use the last six layers of BERT as the cross encoder. The cross encoder takes the text embeddings as input and fuses the image embeddings in the cross-attention module at each layer. Similar to the multi-head Cross-Attention (CA) in PE, the text embeddings serve as queries (q), while the image embeddings act as keys (k) and values (v). The [CLS] embedding of features output by the cross encoder, is projected into 2-dimensional space via an ITM head (i.e., an MLP), yielding the predicted image-text matching probability $\hat{p}(I,T)$. The Image-Text Matching loss is defined as:
@@
Image-Text Contrastive Learning. This constrain targets to align image-text pair via contrastive learning [48, 5, 22]. Formally, given the pose-aware image feature $f_{V}=\{f_{cls},f_{pat}\}$, where $cls$ and $pat$ are the CLS embedding and patch embeddings, respectively. To harvest a comprehensive global image representation, we consider both the patch embedding and the CLS embedding:
@@
where $\text{AVG}(f_{pat})$ is the average feature of tokens in $f_{pat}$. Following previous practice [48, 5, 22], we deploy the initial six transformer layers of BERT [10] for text encoding. The text $T$ is tokenized and prefixed with a single [CLS] token before being input into the encoder. This process yields a text embedding $f_{T}$. Mirroring the image encoding, we apply the same procedure to $f_{T}$, combining text and class embeddings as per Eq. 4, to obtain the global text representation $f_{t}$. The image-to-text similarity within the batch is defined as follows:
@@
Mask Language Modeling (MLM) predicts the masked words in the text based on the matched image. We randomly mask [^2] The replacement rules are as follows: $10\%$ of the tokens are replaced with random tokens, $10\%$ remain unchanged, and $80\%$ are replaced with [MASK]. the input tokens with a probability of $25\%$. The person image $I$ and the corresponding masked text $\hat{T}$ pair are then fed into the cross encoder, followed by an MLM head (an MLP with softmax) to predict the masked tokens. We minimize the cross-entropy loss:
@@
Figure 5: Qualitative Results. Examples of top-5 person anomaly search results with text queries for anomaly actions (top) and normal actions (bottom). Matched images are marked by green rectangles, mismatched images are marked in red, and blue boxes indicate cases where the ID matches but the behavior does not. The parts of the queries that describe appearance, action, and background are highlighted in green, red, and orange.
@@ 5 Experiment
Dataset and Evaluation Metrics. The constructed PAB serves as the evaluation benchmark. To evaluate the performance of Text-based Person Anomaly Search, we adopt recall rates R@K and mAP. Given a query text, all test images are ranked according to their matching probability with the query. If the image that perfectly matches the content of the text description (appearance, actions, background, etc.) is among the top k images, the search is considered successful. Mean Average Precision (mAP) refers to the average area under the precision-recall curve across all queries. Our experimental results report R@1, R@5, R@10, and mAP. Higher recall rates and mAP indicates better results.
Implementation Details. We train the Cross-Modal Pose-aware (CMP) model for 30 epochs with mini-batch size as $22$. We adopt AdamW  [30] optimizer with a weight decay of $0.01$. The learning rate linearly decreases from $1\times 10^{-4}$ to $1\times 10^{-5}$. CMP is initialized with the state-of-the-art pre-trained model for Text-based Person Search [48]. CMP has $230.8$ million trainable parameters, with $86.8M$, $66.4M$, and $59.1M$ parameters for the image, text, and cross encoders, respectively.
5.1 Quantitative and Qualitative Results.
@@
Identity Match (Traditional) 90.95 98.89 99.44 83.45
Behavior Match (Ours) 79.53 97.93 98.84 88.17
Table 3: Results of the proposed CMP on the proposed dataset on different evaluation settings.
@@
Quantitative Results. We adopt the image-text alignment part of APTM model [48], which is the state-of-the-art method for text-based person retrieval, to form a strong baseline. In Figure 4 (a), Zero-Shot refers to evaluating the performance of Text-based Person Anomaly Search directly using the APTM model pre-trained on a large-scale cross-modal pedestrian dataset. The Baseline is trained on $0.1M$ image-text pairs from PAB and achieves $69.92\%$ R@1, significantly higher than the Zero-Shot performance of $9.40\%$. This result indicates that the synthetic training data of PAB facilitates fine-grained behavior retrieval on real-world test sets. Compared to Baseline, our method employs Identity-based Hard Negative Mining (IHNM) and Pose-aware Image Encoder (PE). After being trained on the same $0.1M$ data, our method achieves an R@1 of $72.80\%$, representing a $2.88\%$ improvement over the Baseline model. If all PAB data ($1M$) is adopted to train our CMP model, the recall@1 is $79.53\%$, improves $6.73\%$ compared to $0.1M$ training images. Similar trends are observed in other evaluation metrics, i.e. R@5, R@10, and mAP. These results show the effectiveness of our proposed dataset and the Cross-Modal Pose-aware approach.
Person Anomaly Search Poses Greater Challenges. Text-based person anomaly search demands a finer-grained understanding of both the pedestrian’s appearance and behavior, thus posing more challenges. Table 3 compares the traditional identity match (the images with the same ID are all treated as the correct matches) and the focused person anomaly search. Our model achieves $90.95\%$ R@1 on the task of traditional Text-based Person Search, which is $11.48\%$ higher than our proposed anomaly search. This is because, compared to pedestrian search, which only requires appearance matching, anomaly retrieval is a more challenging task that demands the model to distinguish fine-grained behavioral information. This task characteristic is critical for security event tracing and locating, emergency response, and many other security applications.
@@
Qualitative Results. We qualitatively evaluate our method on the task of Text-based Person Anomaly Search. Figure 5 shows six fine-grained pedestrian behavior retrieval results via the trained Cross-Modal Pose-aware (CMP) model, with queries for anomalous behavior (left) and normal behavior (right). For each query, we display the top five retrieved images. Correct retrieval results are marked with green boxes. Blue and red boxes indicate incorrect matches. Note that blue boxes denote images that belong to the same ID as the query but do not match the action. To further illustrate the retrieval effects, we highlight the parts of the text queries that describe appearance in green, action in red, and the background in orange. It can be observed that, in addition to appearance and background information, our model can effectively distinguish fine-grained action information. Even the incorrectly matched images displayed in Figure 5 still show some relevance to the query sentences.
5.2 Ablation Studies and Further Discussion
@@
Effectiveness of the key component of CMP. We conduct ablation studies on the key components of the Cross-Modal Pose-aware (CMP) model to show their effectiveness. As shown in rows two to five of the Figure 4(a), for fairness, all model variants are trained for $30$ epochs on $100,000$ image-text pairs. First, we evaluate the effectiveness of Identity-based Hard Negative Mining (IHNM) by comparing the Baseline and M1 models. The results show that adding IHNM improves retrieval performance, indicating the necessity of identity-based hard negative samples. Furthermore, M2 incorporating PE obtains $1.87\%$ R@1 improvement compared to Baseline. It indicates that Pose-aware Image Encoder (PE) is a critical component of CMP. Our CMP method, which utilizes both IHNM and PE, shows significant improvements over the Baseline, M1, and M2 models. Specifically, it achieves an R@1 of $72.80\%$, representing a substantial enhancement in retrieval performance.
Impact of Training Data Scale. In general Vision and Language cross-modal tasks and Text-based Pedestrian Research tasks, the scale of the training dataset typically plays a crucial role. More training data can better meet the data requirements of deep models, leading to improved performance. To thoroughly investigate the effectiveness of PAB, we further explore the impact of data scale during training. Specifically, we train for $30$ epochs on $0\%,10\%,25\%,50\%,\text{and }100\%$ of the synthetic training image-text pairs from PAB, separately. Then the performance on the real-world test set is reported as shown in Figure 4(b). The recall tends to improve as the data scale increases. There is a noticeable performance gain from 00 to $10\%$ and the rate of performance improvement gradually slows down from $10\%$ to $100\%$.

### v2

@@ 3.3 Dataset Analysis
Following the steps above, we have successfully created the Pedestrian Anomaly Behavior (PAB) dataset, a large-scale, richly annotated collection. Its properties and examples are depicted in Figure 2. We compare PAB with other prominent Text-based Person Re-Identification datasets (i.e., CUHK-PEDES [24], ICFG-PEDES [11], and RSTPReid [59]) and Video Anomaly Detection datasets (such as ShanghaiTech [32] and UBnormal [1]) in Table 1, focusing on image count, text annotation quantity and length, data sources, and annotation types. PAB features the following characteristics: (1) A Large Number of Anomalous Behaviors: Unlike general pedestrian datasets that focus on the viewpoint and occlusion of pedestrian images, PAB emphasizes providing a large number of images and textual descriptions of pedestrian anomaly. This complements ordinary Person Re-Identification tasks and presents new challenges. (2) High-Fidelity Images: Compared to pedestrian images from surveillance cameras, which often suffer from poor lighting, blurry textures, the images in PAB are of higher quality due to the generation method we adopted. The synthesized images are reasonable, realistic, and aesthetically pleasing. (3) Specific Textual Descriptions: Relative to existing cross-modal pedestrian datasets, the texts in PAB are longer and provide more detailed information, including the appearance, clothing, actions, and background context of individuals. For pedestrian anomaly retrieval, information about the appearance and actions of individuals is crucial. (4) Diversity: PAB encompasses a wide range of images that vary in terms of appearance, posture, viewpoint, background, and occlusion. Additionally, the text generation approach ensures that the image captions in PAB are sufficiently diverse, too. (5) Large-Scale Image-Text Pairs: PAB comprises over one million image-text pairs, supporting deep cross-modal models in learning better uni-modal features and more inter-modal associations from the data. (6) Rich Annotations: Each image-text pair in PAB is annotated with corresponding action, anomaly, and scene categories, providing additional data information besides person appearance. (7) Less Privacy Concerns: Apart from a small amount of data in the test set sourced from existing public datasets, the vast majority of the data in PAB is synthesized, reducing ethical and legal issues.
@@ 4.1 Pose-aware Image Encoder (PE)
Considering that normal and anomaly actions notably differ in human pose, we devise a pose-aware image encoder that explicitly incorporates human pose to enhance behavior comprehension. For a given image $I$, a human key point detector [7] is utilized to obtain the pose map $P$. Given $P$, we extract pose-aware feature via the Pose Conv module, and then feed the feature into the image encoder to obtain $f_{P}$, while $I$ is directly input into the image encoder to extract image embedding $f_{I}$. The Pose Conv module comprises convolutional layers, batch normalization, and ReLU activation functions to align the input domain.
Image Encoder. Without loss of generality, we deploy Swin Transformer (Swin-B) [29] as the Image Encoder. The input image is initially divided into non-overlapping patches, which are then linearly embedded. These embedded patches are subsequently processed by transformer blocks, comprising Multi-Head Self-Attention and Feed Forward modules, to generate patch embeddings. Each patch embedding encapsulates the information of its corresponding patch. To aggregate the information from all patches, we compute the average of their features, referred to as the [CLS] embedding, and prepend it to the sequence.
Pose-aware Cross-Attention. Pose representation $f_{P}$, after being regularized by Layer Norm, is integrated into the image representation through a multi-head cross-attention module. The output ($f_{CA}$) of the cross-attention can be defined by the following equation:
$f_{CA}$ $=Softmax(\frac{{\bf qk}^{T}}{\sqrt{d}}){\bf v},$ (1)
$s.t.\quad{\bf q}$ $=W_{q}f_{P},\quad{\bf k}=W_{k}f_{I},\quad{\bf v}=W_{v}f_{I}.$ (2)
$q,k,v$ are the query, key, and values matrices of the attention operation respectively, $d$ is the dimension of $k$, and $W_{q},W_{k},W_{v}$ are the weights of the linear layers. Then, we combine the output of CA with the image embedding as the pose-aware representation $f_{V}=f_{I}+f_{CA}$.
@@ 4.2 Identity-based Hard Negative Mining (IHNM)
The key to solving pedestrian anomaly research is effectively establishing the relationship between pedestrian descriptions (with normal or anomaly action) and images. Ideally, for each text, the model should be able to search for an image that matches the appearance, actions, and background described, particularly the action description. Training such a cross-modal search model requires large-scale matching image-text pairs, negative pairs, and hard negative pairs, which are crucial for enabling the model to learn more discriminative features. Based on the construction process of PAB, we propose an Identity-based hard negative pair mining (IHNM) method that provides corresponding hard negative samples for each image and text in the training set. The sampling strategy is illustrated in Figure 3 (c).
Consider a training pair $(I,T)$, where $I$ is generated from anomaly/normal caption $C\in\{C_{a},C_{a+},C_{n}\}$, and $T$ is its re-captioned text. As the normal/anomaly counterpart $\tilde{C}$ of $C$ (both describing the same person) is known according to the OOPS! dataset, we can pick $\tilde{I}$ generated from $\tilde{C}$, and its re-captioned text $\tilde{T}$ can also be chosen accordingly. As shown in the Figure 3 (c), $C$ and $\tilde{C}$ describe the anomalous and normal behaviors of the same pedestrian, respectively. Consequently, the generated images have similar pedestrian appearances and backgrounds but different actions. Intuitively, the image $I$ and the text $\tilde{T}$ form an ID-based hard negative pair. The establishment of hard negative pairs with similar appearances and backgrounds but different actions is critical for the model to learn features that are more discriminative of actions.
@@ 4.3 Cross-modal Modeling
IHNM-Enhanced Image-Text Matching. Image-text matching determines whether an (image, text) pair is matched or not. Typically, well-aligned pairs are considered positive samples. Rather than randomly selecting trivial negatives, we employ our IHNM strategy to identify hard negative pairs. These pairs differ only in the action with its positive counterpart, requiring the model to discern subtle action differences, a capability crucial for the task of person anomaly search. Given an image-text pair, their multi-modal feature is encoded by feeding their representation into a cross encoder. Specifically, we use the last six layers of BERT as the cross encoder. The cross encoder takes the text embeddings as input and fuses the image embeddings in the cross-attention module at each layer. Similar to the multi-head cross-attention in PE, the text embeddings serve as queries (q), while the image embeddings act as keys (k) and values (v). The [CLS] embedding of features output by the cross encoder, is projected into 2-dimensional space via an ITM head (i.e., an MLP), yielding the predicted image-text matching probability $\hat{p}(I,T)$. The Image-Text Matching loss is defined as:
@@
Image-Text Contrastive Learning. This constrain targets to align image-text pair via contrastive learning [51, 5, 22]. Given the pose-aware image feature $f_{V}$, we split it to $\{f_{cls},f_{pat}\}$, where $cls$ and $pat$ are the CLS embedding and patch embeddings, respectively. To obtain a comprehensive global image representation, we consider both the patch embedding and the CLS embedding:
@@
where $\text{AVG}(f_{pat})$ is the average feature of tokens in $f_{pat}$. Following previous practice [51, 5], we deploy the initial six transformer layers of BERT [10] for text encoding. The text $T$ is tokenized and prefixed with a single [CLS] token before being input into the encoder. This process yields a text embedding $f_{T}$. Mirroring the image encoding, we apply the same procedure to $f_{T}$, combining text and class embeddings as per Eq. 4, to obtain the global text representation $f_{t}$. The image-to-text similarity within the batch is defined as follows:
@@
Mask Language Modeling (MLM) predicts the masked words in the text based on the matched image. We randomly enable the masking strategy [^2] The masking strategy is: $10\%$ of the tokens are replaced with random tokens, $10\%$ remain unchanged, and $80\%$ are replaced with [MASK]. with a probability of $25\%$. The person image $I$ and the corresponding masked text $\hat{T}$ pair are then fed into the cross encoder, followed by an MLM head (an MLP with softmax) to predict the masked tokens. We minimize the cross-entropy loss:
@@ 5 Experiment
Method #Data R@1 R@5 R@10 mAP
RaSa [5] - 21.74 27.30 27.96 24.35
APTM [51] - 22.90 45.80 52.38 33.56
IRRA [22] - 30.59 59.61 68.91 44.41
CLIP [37] - 47.57 81.55 89.03 62.73
X-VLM [55] - 71.94 97.78 98.99 83.96
APTM [51] 0.1M 72.14 95.30 97.17 82.78
IRRA [22] 0.1M 76.39 97.62 99.14 86.33
CLIP [37] 0.1M 77.60 98.84 99.75 87.35
RaSa [5] 0.1M 80.79 98.89 99.65 89.20
X-VLM [55] 0.1M 81.95 98.84 99.19 89.86
CMP (Ours) 0.1M 83.06 98.89 99.49 90.41
CMP (Ours) 1M 84.93 99.09 99.75 91.66
Table 2: Quantitative results of our proposed method and compared methods on the proposed dataset PAB. The best result is indicated in bold, while the second best is underlined.
Method R@1 R@5 R@10 mAP
APTM [51] 27.86 40.41 46.77 22.61
IRRA [22] 40.28 57.24 65.98 33.53
CLIP [37] 51.60 68.31 76.43 43.05
X-VLM [55] 52.33 66.73 72.54 40.87
RaSa [5] 54.12 70.32 75.96 39.71
CMP 54.12 71.07 77.90 43.13
CMP (1M) 55.23 71.67 77.99 44.35
Table 3: Comparisons with existing methods in OOD setting. The unseen test set UCC is extracted from the UCF-Crime [43] dataset.
Figure 4: Qualitative Results. Top-5 anomaly search results for text queries: anomaly actions (top) and normal actions (bottom). Green rectangles indicate correct matches, red for mismatches, and blue for ID matches with behavior mismatches. Query parts (appearance, action, background) are highlighted in green, red, and orange, respectively.
Dataset and Evaluation Metrics. The constructed PAB serves as the evaluation benchmark. To evaluate the performance of text-based person anomaly search, we adopt recall rates R@K and mAP. Given a query text, all test images are ranked according to their matching probability with the query. If the image that perfectly matches the content of the text description (appearance, actions, background, etc.) is among the top k images, the search is considered successful. Mean Average Precision (mAP) refers to the average area under the precision-recall curve across all queries. We report R@1, R@5, R@10, and mAP. Higher recall rates and mAP indicates better results.
Implementation Details. We train the Cross-Modal Pose-aware (CMP) model for 30 epochs with mini-batch size of $22$. We adopt AdamW  [30] optimizer with a weight decay of $0.01$. The learning rate linearly decreases from $1\times 10^{-4}$ to $1\times 10^{-5}$. CMP has $230.8$ million trainable parameters, with $86.8M$, $66.4M$, and $59.1M$ parameters for the image, text, and cross encoders, respectively. The encoder weights are initialized by X-VLM [55].
@@
Identity Search (Traditional) 94.34 99.39 99.85 88.01
Behavior Search (Ours) 84.93 99.09 99.75 91.66
Table 4: Comparison of different evaluation settings with CMP.
Method PE IHNM R@1 R@5 R@10 mAP
Baseline 83.47 99.04 99.80 90.94
M1 $\checkmark$ 84.13 99.14 99.65 91.20
M2 $\checkmark$ 84.48 98.94 99.60 91.36
CMP (Ours) $\checkmark$ $\checkmark$ 84.93 99.09 99.75 91.66
Table 5: Ablation studies on the key component of our proposed method, i.e., Pose-aware Image Encoder (PE) and Identity-based Hard Negative Mining (IHNM).
@@
Quantitative Results. In Table 2, we compare our method with a wide range of possible solutions, including three state-of-the-art text-based person search methods (APTM [51], IRRA [22], and RaSa [5]) and two state-of-the-art vision-language pre-training models (CLIP [37] and X-VLM [55]). Firstly, X-VLM [55] achieves the highest zero-shot results, with $71.94\%$ R@1, $97.78\%$ R@5, $98.99\%$ R@10, and $83.96\%$ mAP, indicating that PAB presents a significant challenge. After training 30 epochs on the same $0.1M$ image-text pairs from PAB, X-VLM [55] achieves $81.95\%$ R@1, significantly higher than the other four methods. Therefore, we adopt the X-VLM [55] model to form a strong baseline. Compared to Baseline, our Cross-Modal Pose-aware (CMP) method employs Pose-aware Image Encoder (PE) and Identity-based Hard Negative Mining (IHNM) and achieves an R@1 of $83.06\%$, representing a $+1.11\%$ increase over the Baseline model. If all PAB data ($1M$) is adopted to train our CMP model, the recall@1 is $84.93\%$, improving $+1.87\%$ compared to $0.1M$ training images. Similar trends are observed in other evaluation metrics, i.e., R@5, R@10, and mAP. These results verify the effectiveness of CMP.
OOD Setting. To evaluate the scalability of the CMP model, we extract a new real-world test set from UCF-Crime [43] for an Out-of-Distribution (OOD) test, and we call it UCC. Specifically, we select keyframes from $13$ types of abnormal and normal videos in UCF-Crime, leverage Qwen2-VL [4] to generate text descriptions for these video frames. It results $5,320$ image-text pairs as the OOD test set. As shown in Table 3, our model CMP achieves superior performance, $54.12\%$ R@1 and $43.13\%$ mAP, compared to other state-of-the-art methods (also trained on $0.1M$ PAB data). After training with complete $1M$ PAB training data, CMP further improves by $1.11\%$ in R@1. These experimental results show that our model exhibits competitive adaptability when facing unseen datasets.
Behavior Search vs. Identity Search. Text-based person anomaly search poses significantly greater challenges than traditional identity-based person search, as it requires a fine-grained understanding of both appearance and behavior. As shown in Table 4, while traditional methods treat all images with the same ID as correct matches, our anomaly search task demands precise localization of specific behaviors. This distinction is evident in our experimental results: our model achieves 94.34% R@1 on traditional text-based person search, outperforming its performance on anomaly search by 9.41%. The performance gap stems from the fundamental difference between the two tasks: identity search primarily relies on appearance, whereas anomaly retrieval necessitates distinguishing subtle behavioral patterns. Such capability is crucial for real-world applications, including security event tracing, emergency response, and other scenarios where requires precise behavior localization.
@@
Qualitative Results. We qualitatively evaluate our method on the task of Text-based Person Anomaly Search. Figure 4 shows six fine-grained pedestrian behavior retrieval results via the trained Cross-Modal Pose-aware (CMP) model, with queries for anomalous behavior (top) and normal behavior (bottom). For each query, we display the top five retrieved images. Correct retrieval results are marked with green boxes. Blue and red boxes indicate incorrect matches. Note that blue boxes denote images that belong to the same ID as the query but do not match the action. To further illustrate the retrieval effects, we highlight the parts of the text queries that describe appearance in green, action in red, and the background in orange. It can be observed that, in addition to appearance and background information, our model can effectively distinguish fine-grained action information. Even the incorrectly matched images displayed in Figure 4 still show reasonable relevance to the query sentences.
@@
Effect of the Key Components. We conduct ablation studies on the key components of the Cross-Modal Pose-aware (CMP) model. As shown in Table 5, for fairness, all variants are trained for $30$ epochs on $1M$ image-text pairs. First, we evaluate the effectiveness of Pose-aware Image Encoder (PE) by comparing the Baseline and M1 models. It indicates that Pose-aware Image Encoder (PE) is a critical component of CMP. Furthermore, M2 incorporating Identity-based Hard Negative Mining (IHNM) obtains $1.01\%$ R@1 improvement compared to the Baseline. The results show that adding IHNM improves the ability in discriminating the fine-grained behavior. Our CMP method equipped with both PE and IHNM arrives at the best performance.
Figure 5: Ablation studies on synthetic training data scale. We gradually increase PAB training data from 0% to 100% .
Effect of Training Data. We further study the impact of training data scale on model performance in Fig. 5. Considering the hard negative sampling, our method with 10% training data has already achieves competitive recall rate.